OpenStudy (anonymous):

Can someone help me set up the formula? d=rt (d=the distance, r=the average speed, and t=the elaspsed time) A missile travels at the rate of 6,000 mi/hr. Find the distance if it were to travel 15 minutes.

OpenStudy (anonymous):

6000 mi/hr=6000 mi/(60 mins)=6000 mi/4*15 mins=6000/4 (mi/15 mins)

OpenStudy (anonymous):

Can you write that different? I am having a hard time following, sorry.

OpenStudy (anonymous):

6000 mi/hr=6000 mi/(60 mins) get this part?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

6000 mi/(60 mins)=6000 mi/4*15 mins this?

OpenStudy (anonymous):

6000 mi/4*15 mins=6000/4 (mi/15 mins) finally?